OpenStudy (anonymous):

What is the Derivative of t^(3/2) (2+t^(1/2)) Answer is 3t^(1/2)+2t I can't seem to get it.

OpenStudy (anonymous):

you have two functions of 't', are you using the product?

OpenStudy (anonymous):

or multiply through the brackets and use exponent rules before taking the derivative. Which way have you tried?

OpenStudy (anonymous):

All the ways I guess lol Wait so I multiply t^(3/2) through the parens first?

OpenStudy (anonymous):

yes, you can approach it that way...it leads to two terms, each a straight derivative.

OpenStudy (anonymous):

ok wait what do you get for the first step 2t^(3/2)+t^(4/2)?

OpenStudy (anonymous):

Damn I got it now! THanks soo much

OpenStudy (anonymous):

I thought we are supposed to do whats inside the parens first?

OpenStudy (anonymous):

any algebraic steps you do to the function don't alter it in any way, so sometimes its just a matter of massaging into into the most convenient form

OpenStudy (anonymous):

alternately, if you were to go directly from the form given: \[\huge f(t)=t^{\frac{3}{2}}(2+t^{\frac{1}{2}})\] \[\huge f'(t)=uv'+u'v\] \[u=t^{\frac{3}{2}}\] \[u' = \frac{3}{2}t^{\frac{1}{2}}\] \[v=2+t^{\frac{1}{2}}\] \[v'=\frac{1}{2}t^{-\frac{1}{2}}\]

OpenStudy (anonymous):

ok thank you